# SWAN: Preprocessing SGD Enables Adam-Level Performance On LLM Training With Significant Memory Reduction

### v1

@@ Appendix G Proof of Proposition 1
Appendix H Proof of Theorem 2
Proof.
Assume that we have a function $f({\mathbf{W}})$ defined on ${\mathbf{W}}\in\mathbb{R}^{m\times n}$ that we want to minimize, and the global minimum is at ${\mathbf{W}}^{*}$ (without loss of generality, let ${\mathbf{W}}^{*}=0$). Then our problem can be written as:
$\min_{{\mathbf{W}}\in\mathbb{R}^{m\times n}}f({\mathbf{W}})$ (22)
Assume $f$ is a twice differentiable function and has a Lipschitz continuous Hessian around ${\mathbf{W}}^{*}$, then both $\mathtt{GradWhitening}$ and Newton’s method approximate the loss equation 22 up to second order (ignoring first-order terms without loss of generality) as:
$\min_{{\mathbf{w}}\in\mathbb{R}^{mn}}\frac{1}{2}{\mathbf{w}}^{T}{\mathbf{H}}{\mathbf{w}},$ (23)
where ${\mathbf{w}}\in\mathbb{R}^{mn\times 1}$ is the flattened version of ${\mathbf{W}}$, and ${\mathbf{H}}\in\mathbb{R}^{mn\times mn}$ is the Hessian matrix of $f$ at ${\mathbf{W}}^{*}$. We may apply second-order Newton’s method to optimize equation 23, which gives the following update:
${\mathbf{w}}^{t+1}={\mathbf{w}}^{t}-\eta{\mathbf{H}}^{-1}\nabla f({\mathbf{w}}^{t}).$ (24)
However, this update is very expensive and requires $\mathcal{O}(m^{3}n^{3})$ computation. SWANcan then be viewed as a fast approximation to estimate the update of equation 24. Under our assumption on Hessian structure, we have:
${\mathbf{H}}=\text{diag}(\tilde{{\mathbf{H}}},\tilde{{\mathbf{H}}},\dots,\tilde{{\mathbf{H}}},\dots,\tilde{{\mathbf{H}}})$ (25)
where $\tilde{{\mathbf{H}}}\in\mathbb{R}^{m\times m}$.
Under this assumption, the optimization problem equation 23 can be rewritten as:
$\min_{{\mathbf{W}}\in\mathbb{R}^{m\times n}}\frac{1}{2}\text{Tr}\left[{\mathbf{W}}^{\top}\tilde{{\mathbf{H}}}{\mathbf{W}}\right],$ (26)
For this model, the gradient of ${\mathbf{W}}$ is given by:
$\nabla f({\mathbf{W}})=\tilde{{\mathbf{H}}}{\mathbf{W}}.$
Let ${\mathbf{W}}$ have rank $m$, and assume $n>m$. We say ${\mathbf{W}}$ is now over-parameterized in the sense that its dimensionality is now larger than the dimensionality of the second-order information $\tilde{{\mathbf{H}}}$. It is now clear that the second-order information is losslessly compressed into gradients: one can exactly recover the Hessian via
${\mathbf{H}}=\text{diag}(\nabla f({\mathbf{W}}){\mathbf{W}}^{-1},\nabla f({\mathbf{W}}){\mathbf{W}}^{-1},\dots,\nabla f({\mathbf{W}}){\mathbf{W}}^{-1},\dots,\nabla f({\mathbf{W}}){\mathbf{W}}^{-1}).$
Given this fact, one can easily estimate Newton’s update ${\mathbf{H}}^{-1}\nabla f({\mathbf{w}}^{t})$ without explicitly estimating ${\mathbf{H}}^{-1}$ (or $\tilde{{\mathbf{H}}}$). Notice that:
${\mathbf{H}}={\mathbf{I}}_{m}\otimes\tilde{{\mathbf{H}}}$
where ${\mathbf{I}}_{m}$ is the $m\times m$ identity matrix and $\otimes$ denotes the Kronecker product. Given that
${\mathbf{H}}^{-1}={\mathbf{I}}_{m}\otimes\tilde{{\mathbf{H}}}^{-1},$
we have
${\mathbf{H}}^{-1}\nabla f({\mathbf{w}}^{t})=({\mathbf{I}}_{m}\otimes\tilde{{\mathbf{H}}}^{-1})\,\text{vec}(\nabla f({\mathbf{W}}^{t}))=\text{vec}\left(\tilde{{\mathbf{H}}}^{-1}\nabla f({\mathbf{W}}^{t})\right).$
Reshaping back to matrix form, we can rewrite the update rule as:
$\tilde{{\mathbf{H}}}^{-1}\nabla f({\mathbf{W}}^{t}).$
Next, since $\tilde{{\mathbf{H}}}$ is positive definite:
$\nabla f({\mathbf{W}}^{t})=\tilde{{\mathbf{H}}}{\mathbf{W}}^{t}={\mathbf{Q}}\Lambda{\mathbf{Q}}^{\top}{\mathbf{W}}^{t},$
where ${\mathbf{Q}}\Lambda{\mathbf{Q}}^{\top}$ is the eigen decomposition of $\tilde{{\mathbf{H}}}$. Assume further that ${\mathbf{W}}^{t}$ is initialized as orthogonal; one can immediately recognize that ${\mathbf{Q}}$ is also the left singular vector matrix of $\nabla f({\mathbf{W}}^{t})$. Therefore, the matrix inversion of $\tilde{{\mathbf{H}}}$ required in Newton’s update equation 24 can be easily achieved by simply deleting the $\Lambda$ term. That is,
$\tilde{{\mathbf{H}}}^{-1}\nabla f({\mathbf{W}}^{t})={\mathbf{Q}}\Lambda^{-1}{\mathbf{Q}}^{\top}{\mathbf{Q}}\Lambda{\mathbf{Q}}^{\top}{\mathbf{W}}^{t}$
$=$ ${\mathbf{Q}}\left[{\mathbf{Q}}^{\top}{\mathbf{W}}^{t}\right]\overset{{\mathbf{U}}:={\mathbf{Q}},{\mathbf{V}}:=({\mathbf{W}}^{t})^{\top}{\mathbf{Q}}}{=}{\mathbf{U}}{\mathbf{V}}^{\top}=\mathcal{P}\left(\nabla f({\mathbf{W}}^{t})\right)={\mathbf{G}}_{\text{whitened}}.$
This is exactly the $\mathtt{GradWhitening}$-modified gradient update. ∎
Appendix I Implementation details
@@
We describe the implementation setups for SWAN used in LLM pre-training tasks. To enable a more straightforward and comparable analysis, we simply replicate the setting of Zhao et al. (2024a), under exactly the same model configs and optimizer configs. This include the same model architecture, tokenizer, batch size, context length, learning rate scheduler, etc. The only exception is that since SWAN does not need warm up, we use zero-warm up ratio for SWAN. But we still ensure that the learning rates of all methods are fully decayed to the same target learning rate (0.1 $\times$ inital learning rate). Across all model sizes, we use the same hyperparameter settings and the same learning rate tuning procedure. The only exception is that for model of size 1.3B: as we already know that larger model requires smaller learning rates, we conduct learning search for Adam and Galore over a smaller but more fine grained grid of $\{0.001,0.0007,0.0005,0.0003,0.0001\}$. As a result, the optimal learning rate found for Adam on 1.3B is 0.0007.
@@
Since SWAN utilizes matrix-level operations on gradients, it can only be applied on 2D parameters. Therefore, in our experiments we only apply $SWAN{}$ on all linear projection weights in transformer blocks. Similar to Galore (Zhao et al., 2024a), the rest of the non-Linear parameters still uses Adam as default choice. Therefore, we follow the learning rate setup of Galore, where we fix a global learning rate 0.001 across all model sizes and all modules. Then, for the linear projection modules where SWAN is applied, we simply apply a scaling factor $\alpha$ on top of the global learning rate. We choose $\alpha=1$ for all model sizes except for 1.3 B case, where we used $\alpha=0.3$, resulting an effective learning rate of $0.0003$. Note that this choice is quite random; and it is not extensively grid-searched.
Applying $\mathtt{GradWhitening}$
We found that one strategy that lead to better convergence is to first convert the gradient matrice to FP32, perform $\mathtt{GradWhitening}$, and then convert back to BF16. This is the default setup in Section 5.3. In practice, we found that using a step size $\beta\neq 0.5$ in the $\mathtt{GradWhitening}$ operator (Algorithm 2) can improve its convergence. We found that setting $\beta=0.8$ with iteration number $k=10$ generally give satisfactory performance. This can be further optimized and we leave it for future work. Alternatively, $\mathtt{GradWhitening}$ can also be done in BF16 precision. The setup we use for BF16 precision is $\beta=0.85$, $k=6$. These hyperparameters are selected on randomly sampled matrices via minimizing inverse square root estimation loss; then directly applied to LLM without any further tuning. We found that this usually translates well.

### v2

@@ Appendix G Proof of Proposition 1
Appendix H Implementation details
@@
We describe the implementation setups for SWAN used in LLM pre-training tasks. To enable a more straightforward and comparable analysis, we simply replicate the setting of Zhao et al. (2024a), under exactly the same model configs and optimizer configs. This includes the same model architecture, tokenizer, batch size, context length, learning rate scheduler, etc. The only exception is that since SWAN does not need warm-up, we use zero-warm-up ratio for SWAN. But we still ensure that the learning rates of all methods are fully decayed to the same target learning rate (0.1 $\times$ initial learning rate). Across all model sizes, we use the same hyperparameter settings and the same learning rate tuning procedure. The only exception is that for a model of size 1.3B: as we already know that a larger model requires smaller learning rates, we conduct a learning search for Adam over a smaller but more fine-grained grid of $\{0.001,0.0007,0.0005,0.0003,0.0001\}$. As a result, the optimal learning rate found for Adam on 1.3B is 0.0007.
@@
Since SWAN utilizes matrix-level operations on gradients, it can only be applied to 2D parameters. Therefore, in our experiments, we only apply $SWAN{}$ on all linear projection weights in transformer blocks. Similar to Galore (Zhao et al., 2024a), the rest of the non-linear parameters still uses Adam as the default choice. Therefore, we follow the learning rate setup of Galore, where we fix a global learning rate 0.001 across all model sizes and all modules. Then, for the linear projection modules where SWAN is applied, we simply apply a scaling factor $\alpha$ on top of the global learning rate. We choose $\alpha=1$ for all model sizes except for 1.3 B case. This is because we observe that the optimal learning rate of Adam under 1.3B becomes smaller than 0.001, hence we also reduce the learning rate on SWAN, where we used $\alpha=0.3$, resulting an effective learning rate of $0.0003$. Note that this choice is quite arbitrary, and it is not extensively grid-searched. This is to show how the method work out-of-the box, without extensive problem-specific tuning.
Implementation of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$
For $\mathtt{GradNorm}$, we simply set $\bar{{\bm{g}}}=0$ to simplify computation, as we found in practice this often has small scales. For $\mathtt{GradWhitening}$, before N-S iteration, we further normalize its input matrix by its Frobenius norm. This is applied in all our ablation studies as well, regardless of whether the gradient is processed by $\mathtt{GradNorm}$. For the N-S iteration, we found that using a step size $\beta\neq 0.5$ in the $\mathtt{GradWhitening}$ operator (Algorithm 2) can improve its convergence. We set $\beta=0.8$ with iteration number $k=10$ and it generally gives satisfactory performance. $\mathtt{GradWhitening}$ can be performed in both FP32 and BF16 format. In FP32 format, we first convert the gradient matrice to FP32, perform $\mathtt{GradWhitening}$, and then convert back to BF16. This is the default setup in Section 6.1 unless stated otherwise. Alternatively, $\mathtt{GradWhitening}$ can also be done in BF16 precision. The setup we use for BF16 precision is $\beta=0.85$, $k=6$. These hyperparameters are selected on randomly sampled matrices via minimizing inverse square root estimation loss; then directly applied to LLM without any further tuning. We found that this usually translates well. Note that this might be further optimized similar to (Jordan et al., 2024), we leave it to future work as N-S performance optimization is orthogonal to our key contribution.